OpenStudy (anonymous):

Find the remaining trigonometric ratios. tan(a)=2, 0

OpenStudy (blockcolder):

Draw a right triangle.

OpenStudy (anonymous):

Then..?

OpenStudy (blockcolder):

Use SOHCAHTOA to find sin, cos. Use the fact that sec=1/cos, csc=1/sin, cot=1/tan to get the rest.

OpenStudy (anonymous):

 What would the hyp. be?

OpenStudy (blockcolder):

Use the Pythagorean Theorem.

OpenStudy (blockcolder):

BTW, the 2 and the 1 are in the wrong spots, if the angle on the right side is a.